OpenStudy (anonymous):

You were named District Manager of the Month at BeBop Records due to a hiring study you conducted. Your study showed that each of 30 salespersons that you supervise average $50,000 in sales each month. Your study also showed that for each new salesperson hired, the average monthly sales would decrease $1,000. You presented this study to the company's CEO. What is the maximum average sales the company can expect with all it's salespersons? Hint: Let x represent the number of new salespeople. Total monthly sales = (number of salespersons)(average monthly sales each)

OpenStudy (anonymous):

Total monthly sales = (number of salespersons)(average monthly sales each) = (30 + x)(50,000 - x)

OpenStudy (anonymous):

@garrettlung

OpenStudy (anonymous):

it also comes with answer choices did you want me to show them ?

OpenStudy (anonymous):

i'm not so sure about this...

OpenStudy (michele_laino):

I'm very sorry, I don't know your answer since I'm not good with financial mathematics

OpenStudy (anonymous):

The product of consecutive odd integers is 575. What is the largest integer? here's the other one I'm stuck on

OpenStudy (michele_laino):

here we can write this equation: \[\Large \left( {2n - 1} \right)\left( {2n + 1} \right) = 575\] since 2n-1 and 2n+1 are two consecutive odd numbers

OpenStudy (michele_laino):

so, after a simplification at left side, we get: \[\Large 4{n^2} - 1 = 575\]

OpenStudy (michele_laino):

please continue

OpenStudy (michele_laino):

hint: you have to solve that equation for n

OpenStudy (anonymous):

do would I use 2 ?

OpenStudy (anonymous):

i got 15

OpenStudy (michele_laino):

hint: \[\Large {n^2} = \frac{{576}}{4}\]

OpenStudy (michele_laino):

what is n?

OpenStudy (anonymous):

I'm still working out the problem

OpenStudy (anonymous):

144

OpenStudy (anonymous):

wait sorry

OpenStudy (anonymous):

143+3/4

OpenStudy (michele_laino):

n^2=144, and n?

OpenStudy (anonymous):

sorry i got it n=12

OpenStudy (michele_laino):

that's right! so, what is: 2*n+1=...?

OpenStudy (anonymous):

25

OpenStudy (michele_laino):

correct! the 2 numbers are: 23, and 25

OpenStudy (anonymous):

so do I put in 25 ?

OpenStudy (michele_laino):

yes! 25

OpenStudy (anonymous):

okay